Analyze and solve the following problems.

a. What is the probability that when you roll a die, the outcome is an even number?

b. What is the probability that when you determine the blood type of an individual his or her blood type is type O?

c. In a standard deck of cards, what is the probability of drawing a card that is a heart or an ace?

d. A coin is tossed thrice. What is the probability of obtaining exactly 2 heads occur?

e. According to the recent news, the probability of raining today is 65%. What is the probability of not raining today?


1
Expert's answer
2022-04-05T10:02:16-0400

a There 6 possible results and 3 of them are successful. So probability is 3/6=1/2

b. There are 4 blood types. 1 of them is successful. So probability is 1/4

c. There are 52 cards in deck. 13 hearts + 3 aces are successful. So, probability is 16/52=4/13

d. Probability of head is 0.5, probability of not not head is 0.5.

It can be 3 ways of heads from 3 tosses. So, probably is 3*( 0.5)3=0.375

e. Probability of not raining today is 100-65=35%